# Periodic robust robotic rock chop via virtual model control

### v1

@@ III-B Virtual model control for cutting
Finally, to ensure the knife moves primarily within the plane that defines the slice, we introduce four additional points $\boldsymbol{p}_{3},\dots,\boldsymbol{p}_{6}$ on the knife (Fig. 3(d)) and connect each to a virtual mass free to slide along the desired slicing plane, whose position is denoted by the reference $\boldsymbol{r}_{S}$ Each pair is linked by its own spring and damper. This virtual mechanism enforces a plane-constrained motion, yielding a cleaner cut. To produce multiple slices in sequence, once a slice is completed and the knife has returned to its raised position, we shift the slicing plane $\boldsymbol{r}_{S}$, $\boldsymbol{p}_{a}$, $\boldsymbol{r}_{2,1}$, and $\boldsymbol{r}_{2,2}$ to the next desired location.
IV Experimental platform and cutting motion
IV-A Robot platform and control architecture
@@
The virtual model controller outlined in the previous section is implemented through the VMRobotControl.jl package [45], a platform-independent framework designed for virtual model control in robotic systems. This package enables the construction and simulation of both the robot and virtual components. Given the robot’s URDF representation, the package offers several primitives to define the virtual model controller, to compute its dynamics, and to generate the torque commands for each joint.
@@
Our experimental platform consists of a 7-DoF Franka Research 3 arm, with a maximum payload of 3 kg. To collect force data, a ROBOTIQ FT 300-S force-torque sensor is mounted at the robot wrist, providing force and torque measurements at 100 Hz with a noise level of approximately 0.1 N. A knife is rigidly connected to the sensor using a 3D-printed attachment as shown in Fig. 4. Thanks to the platform independence of the virtual model design, the cutting mechanism was also implemented on a humanoid robot. Sciurus17, a lightweight and low-cost platform developed by RT Corp. Scirus17 features 17 degrees of freedom (DoF) distributed across its upper body (with 7 DoF per arm, 2 DoF for the head and 1 DoF for the torso). The robot utilizes Dynamixel motors with current control, and the conversion from torque to current is approximated by means of a linear relation based on the performance characteristics of the motor.
Figure 4: Experiment setup: robot arm with force sensor and knife rigidly attached.
@@
The rocking motion was initially validated via simulations. However, because of the computational cost and potential inaccuracies associated with simulating contact forces for cutting, further tuning was performed during real-world experiments. The controller runs on a standard computer. The overall control architecture is depicted in Fig. 5. Virtual forces $\boldsymbol{f}$ generated by the virtual mechanisms are computed and translated into joint torques $\boldsymbol{\tau}$ using (1). The computation of the current state of the virtual mechanisms, such as the elongation of virtual springs and the velocities used by the virtual dampers, is based solely on the robot’s joint position $\boldsymbol{q}$ and velocity $\dot{\boldsymbol{q}}$. All force computations and their conversion into joint torques rely on forward kinematics and the associated Jacobians. This represents a significant departure from traditional motion planning, where execution is typically based on inverse kinematics.
@@
IV-B Emergence of cutting motion in experiments
@@
The desired rocking motion is not predetermined but emerges from the coordinated action of the several parts of the virtual mechanism in Fig. 3. The virtual mechanisms constrain the motion of the blade by enforcing (compliant) contact points and ‘soft’ constraints. Switching of the reference at specific points induces a cyclical motion of the knife, which is also modulated by the interaction among robot dynamics, compliant virtual mechanism constraints, and contacts with the environment. A first illustration is provided in Fig. 6(d), whose control parameters are specified in Table I, A-1. Using courgette as the cutting target, the snapshots in Fig. 6(a) illustrate the continuous contact between the knife and the cutting board while the knife rolls back and forth. The successful separation of the courgette is also seen in the snapshots. Since the motion is not pre-programmed, variations in the knife geometry or cutting board features will yield different motions.
Fig. 6(b) shows the position of two springs responsible for the rocking motion. The sliding of the virtual mass ensures that $\boldsymbol{p}_{a}$ stays close to $\boldsymbol{p}_{1}$ and no additional force is applied in the rocking direction. The displacement between $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{a}$ along the $z$ axis pulls the knife towards the cutting board and ensures contact. The knife moves to the next slice position when raised. Fig. 6(c) and 6(d) show the end-effector’s position and force, orientation and torque. The resulting end-effector’s motion is a fairly complicated trajectory, difficult to design manually. One would need to know the geometry of the knife to decide where the end-effector should be while rolling the knife back and forth.
@@
V Energy-based analysis of the periodic motion
Energy considerations provide insight into the emerging periodic motion. The mechanical energy of the robot, $E_{R}$, satisfies the following standard passivity inequality [46, 34, 43]
where $\boldsymbol{q}$ is the vector of joint coordinates of the robot manipulator, $\boldsymbol{\tau}$ are the related motor torques, and $(\boldsymbol{s}_{i},\boldsymbol{d}_{i})$ are additional co-located positions/forces variables, accounting for generic external actions to the robot. The inequality takes into account unavoidable losses due to internal mechanical dissipation.
The controller is also a virtual mechanism, whose mechanical energy satisfies
$W_{\mathrm{dissipation}}\geq 0$ accounts for the power dissipated in the dampers of the virtual mechanism. $W_{\mathrm{elastic}}$ captures the power flow caused by the switching of the references $\boldsymbol{r}_{2,1}$, $\boldsymbol{r}_{2,2}$, and $\boldsymbol{r}_{S}$.
The negative sign of the term $-\sum_{i=1}^{6}\dot{\boldsymbol{p}}_{i}\boldsymbol{f}_{i}$ in (6) is due to the mechanical interconnection between robot and virtual model controller. Through (1), each force $\boldsymbol{f}_{i}$ acting on the robot exerts an equal and opposite force on the virtual mechanism. The total energy of the controlled robot, $E=E_{\mathrm{R}}+E_{\mathrm{VMC}}$, thus satisfies
(7) shows that the energy of the controlled must continuously decrease if no external forces act on the robot (i.e. $\boldsymbol{d}_{i}=0$) and if we do not switch references $\boldsymbol{r}_{2,1}$, $\boldsymbol{r}_{2,2}$, and $\boldsymbol{r}_{S}$ of the virtual mechanism. The robot motion eventually settles in a resting position. For the gravity-compensated robot, this corresponds to a minimum of the elastic potential energy. This is a classical result of passivity-based control [42].
In our case, each time a reference is switched at a specific robot configuration, a bounded amount of energy is either injected into or removed from the system. This action continuously drives the controlled robot into motion. Switching between the cutting and raising phases is governed by different triggers: the transition from raising the knife to cutting is position-based, while the switch from cutting to raising is primarily triggered by the knife’s orientation. However, because the cutting board restricts the motion of the knife, we can also treat the latter as a position-based trigger. When a switch occurs, the spring that was previously connected to the reference $\boldsymbol{r}_{2,1}$ is detached and reattached to $\boldsymbol{r}_{2,2}$, and vice versa. Because of these reference switches, the virtual springs generate new forces, leading to non-zero velocities and thus new dissipation ($W_{\mathrm{dissipation}}>0$) in the virtual dampers. The injected elastic energy is always bounded, due to the position-based switching mechanism.
The periodic cutting motion of the controlled robot results from the balance between the bounded energy injected through reference switching and the energy dissipated by virtual dampers, the internal dissipation of the robot, and the environment. That is, neglecting internal dissipation and environment contributions, for any periodic motions of period $T$ we have, for all $t$,
[…]
The overall system does not accumulate energy indefinitely. The injected energy is either dissipated by the robot’s inherent damping and the virtual dampers, or transferred into the environment (e.g., performing cutting work on the vegetable or contacting the board). Hence, over multiple cycles, the system settles to a bounded regime of energy. Any surplus energy in the system regardless of its source ultimately appears as increased speed at switching or larger cutting forces. Nonetheless, it is dissipated through robot’s physical and virtual damping, or absorbed by subsequent contact with the board or vegetable, preventing unbounded energy growth.
Further insights can be gathered by looking at a toy model, given by a single mass and one degree of freedom. For a single controlled mass we have
$m\ddot{x}+c\dot{x}+k(x-r)=0,$ (9)
where $m>0$ is the mass, $c>0$ is the damping coefficient, and $k>0$ is the spring stiffness. $r$ is the reference, switching between two values, $r\in\{r_{1},r_{2}\}$. $x$ is the mass position. The mechanical energy reads
$E=\frac{1}{2}m\dot{x}^{2}+\frac{1}{2}k(x-r)^{2}$ (10)
from which
$\dot{E}=-c\dot{x}^{2}+W_{r}.$ (11)
$W_{r}$ accounts for the flow of energy due to switching reference. For instance, $W_{r}$ is zero unless a switch occurs. A switch at time $t$ causes an instantaneous (and large!) flow of energy, $W_{r}\neq 0$. The amount of injected energy can be computed by looking at the left and right limits of the switching time $t$, denoted by $t^{-}$ and $t^{+}$. For instance,
${E(t^{+})-E(t^{-})}=\frac{1}{2}k(x(t^{+})-r(t^{+}))^{2}-\frac{1}{2}k(x(t^{-})-r(t^{-})^{2}.$ (12)
If $r\in\{r_{1},r_{2}\}$ and switch times depend on $x$, this injection is always bounded. A periodic motion arises at the balance between injection of energy (switches) and dissipation (during motion). Consider the case that a switch occurs at positions $x_{1}$ and $x_{2}$, as depicted in Fig. 7. This case mirrors the cutting/raising transitions in rocking motion. Formally,
$r$ $=r_{1}\quad\text{if}\quad\bigl{(}\dot{x}\geq 0\,\wedge\,x\leq x_{1}\bigr{)}\;\text{or}\;\bigl{(}\dot{x}\leq 0\,\wedge\,x\leq x_{2}\bigr{)},$
$r$ $=r_{2}\quad\text{if}\quad\bigl{(}\dot{x}\geq 0\,\wedge\,x\geq x_{1}\bigr{)}\;\text{or}\;\bigl{(}\dot{x}\leq 0\,\wedge\,x\geq x_{2}\bigr{)}.$
Figure 7: Switching the reference between $r_{1}$ and $r_{2}$ at positions $x_{1}$ and $x_{2}$. This schematic mirrors the cutting/raising transitions in rocking motion.
By plotting trajectories in the $(x,\dot{x})$ phase plane (Fig. 8), one observes that solutions converge to a closed curve. Intuitively, if the mass moves too fast, the damper dissipates more energy, returning the system toward this closed curve. Conversely, if motion is too slow, less energy is dissipated, and the system gains more momentum until it approaches the same closed curve.
A final analysis can be developed experimentally. Consider a generic cutting cycle identified by the robot switching reference from ‘raising’ to ‘cutting’. The total work done by the robot per cutting cycle can be approximated by the energy
$E_{\mathcal{C}_{j}}\,=\,\sum_{t_{i}\in\mathcal{C}_{j}\setminus\{t_{N}\}}\boldsymbol{F}\bigl{(}t_{i}\bigr{)}\bigl{(}\boldsymbol{x}(t_{i+1})-\boldsymbol{x}(t_{i})\bigr{)},$ (13)
where $\boldsymbol{F}$ is the force read from the force sensor placed at the end-effector while cutting. $\boldsymbol{x}$ is its position. The set $\mathcal{C}_{j}=\{t_{1},\dots,t_{N}\}$ contains the time instants of cycle $j$ at which sample measures are taken. At each cycle, $E_{\mathcal{C}_{j}}$ provides an estimate of the amount of work required to cut, considering all dissipations and interactions (internal, external, food, cutting board).
We start the experiment with the knife initially performing a pure rocking motion with no food. From the 7th to the 16th cycle, a carrot is introduced and cut by the knife, after which the carrot is removed. As illustrated in Fig. 9, the work done for cutting the carrot differs from the work done during baseline rocking motion. However, once the carrot is removed, the system returns to the energy level observed during pure rocking. This behavior shows how the virtual model controller adjusts the robot motion to compensate for external perturbations, and consistently returns to prior energy balance when perturbations are removed.
Figure 8: Phase plane of a mass spring damper system with switching references $r_{1}=1$ and $r_{2}=-1$ and positions $x_{2}=-0.8$ and $x_{1}=0.8$. Trajectories with different initial conditions converge to the same limit cycle.
VI Experimental result and discussion
VI-A Forces, velocities, and frequency of successful cut

### v2

@@ 3.2 Virtual model control for cutting
Finally, to ensure the knife moves primarily within the plane that defines the slice, we introduce four additional points $\boldsymbol{p}_{3},\dots,\boldsymbol{p}_{6}$ on the knife (Fig. 3(d)) and connect each to a virtual mass free to slide along the desired slicing plane, whose position is denoted by the reference $\boldsymbol{r}_{S}$. Each pair is linked by its own spring and damper. This virtual mechanism enforces a plane-constrained motion, yielding a cleaner cut. Attachment points $\boldsymbol{p}_{3}$$\boldsymbol{p}_{6}$ are defined relative to the knife plane with respect to the end effector. These points need not lie on the knife itself; increasing their distance from the blade enhances orientation correction. While three points are theoretically sufficient to constrain orientation, using four provides improved stability without further increasing stiffness. To produce multiple slices in sequence, once a slice is completed and the knife has returned to its raised position, we shift the slicing plane $\boldsymbol{r}_{S}$, $\boldsymbol{p}_{a}$, $\boldsymbol{r}_{2,1}$, and $\boldsymbol{r}_{2,2}$ to the next desired location.
4 Energy-based analysis
Energy considerations provide insight into the emerging periodic motion. The mechanical energy of the robot, $E_{R}$, satisfies the classical passivity inequality Spong (2022); Spong et al. (2005); Ortega et al. (1998)
where $\boldsymbol{q}$ is the vector of joint coordinates of the robot manipulator, $\boldsymbol{\tau}$ are the related motor torques, and $(\boldsymbol{s}_{i},\boldsymbol{d}_{i})$ are additional co-located positions/forces variables, accounting for generic external actions to the robot due to the interaction of the knife with food and cutting board. The inequality takes into account unavoidable losses due to internal mechanical dissipation.
The controller is also a (virtual) mechanism, whose mechanical (virtual) energy satisfies
$W_{\mathrm{dissipation}}\geq 0$ accounts for the power dissipated in the dampers of the virtual mechanism. $W_{\mathrm{elastic}}$ captures the power flow caused by the switching of the references $\boldsymbol{r}_{2,1}$, $\boldsymbol{r}_{2,2}$, and $\boldsymbol{r}_{S}$. The negative sign of the term $\sum_{i=1}^{6}\dot{\boldsymbol{p}}_{i}\boldsymbol{f}_{i}$ in (6) is due to the mechanical interconnection between robot and virtual model controller. Through (1), each force $\boldsymbol{f}_{i}$ acting on the robot exerts an equal and opposite force on the virtual mechanism.
Taking into account the identity $\dot{\boldsymbol{q}}^{T}\boldsymbol{\tau}=\dot{\boldsymbol{q}}^{T}\sum_{i=1}^{6}J_{i}(\boldsymbol{q})^{T}\boldsymbol{f}_{i}=\sum_{i=1}^{6}\dot{\boldsymbol{p}}_{i}^{T}\boldsymbol{f}_{i}$, the total energy of the controlled robot $E=E_{\mathrm{R}}+E_{\mathrm{VMC}}$ satisfies
(7) shows that the energy of the controlled robot must continuously decrease if no external forces act on the robot (i.e. $\boldsymbol{d}_{i}=0$) and if we do not switch references $\boldsymbol{r}_{2,1}$, $\boldsymbol{r}_{2,2}$, and $\boldsymbol{r}_{S}$ of the virtual mechanism. Likewise, the energy necessarily decreases also if the external forces $\boldsymbol{d}_{i}$ represent interaction forces with the food and cutting board. This follows from passivity theory Willems (1972); Ortega et al. (1998); van der Schaft (1999); Secchi et al. (2007), as the complete mechanical system consisting of the controlled robot interacting with food and cutting board can be modelled as the negative feedback interconnection of passive systems. In both scenarios, without switching, the robot motion eventually settles in a resting position. For the gravity-compensated robot, this corresponds to a minimum of the elastic potential energy. This is a classical result of passivity-based control Ortega et al. (2001).
In our case, each time a reference is switched at a specific robot configuration, a bounded amount of energy is injected into or removed from the system. This action continuously drives the controlled robot into motion. The switch between the cutting and raising phases is governed by specific triggers: the transition from raising the knife to cutting is based on the position of the knife, while the switch from cutting to raising is mainly triggered by the orientation of the knife. When a switch occurs, the spring connecting $\boldsymbol{p_{b}}$ to $\boldsymbol{r}_{2,i}$, $i\in\{1,2\}$, is instantly detached from $\boldsymbol{r}_{2,1}$ and attached to $\boldsymbol{r}_{2,2}$ and vice-versa. The injected/removed energy corresponds to the discontinuity in elastic potential energy resulting from the instantaneous change in spring extension.
The periodic cutting motions of the controlled robot, illustrated in the experimental results section below, emerge at the equilibrium between the energy injected via reference switching and the total dissipation from virtual damping, robot internal dissipation, and environmental resistance (i.e. food and cutting board). Along any periodic motions of period $\bar{T}$ the controlled robot must satisfy
$\int_{t}^{t+\bar{T}}\!\!\!\!\!\!\!W_{\mathrm{elastic}}(\tau)d\tau=\!\int_{t}^{t+\bar{T}}\!\!\!\!\!\!\!W_{\mathrm{dissipation}}(\tau)\!+\!\sum_{i}\dot{\boldsymbol{s}}_{i}^{T}(\tau)\boldsymbol{d}_{i}(\tau)d\tau$ (8)
for all times $t$ (where we have neglected the internal dissipation of the robot, for simplicity).
The energy injected at each switch is dissipated by the robot’s inherent damping and the virtual dampers, or transferred to the environment through interaction with the food and the cutting board. Any energy surplus ultimately manifests as increased velocity or higher cutting forces, both of which enhance dissipation. Due to the geometry of the compliant constraints imposed on the knife by the virtual mechanisms, the switch-induced variations primarily affect the knife’s planar velocity and its contact forces. Hence, over multiple cycles, the system settles into a stable periodic regime.
Further insights can be obtained by examining a simplified single-degree-of-freedom model, representing a first-order approximation of the vertical dynamics of $\boldsymbol{p}_{2}$
$m\ddot{z}+c\dot{z}+k(z-r)=0,$ (9)
where $m>0$ is the mass, $c>0$ is the damping coefficient, and $k>0$ is the spring stiffness. $r$ is the reference, switching between two values, $r\in\{r_{1},r_{2}\}$. The mechanical energy reads
$E=\frac{1}{2}m\dot{z}^{2}+\frac{1}{2}k(z-r)^{2}$ (10)
from which
$\dot{E}=-c\dot{z}^{2}+W_{r}.$ (11)
$W_{r}$ accounts for the elastic energy injected or removed at switching times (an impulse). This can be computed by looking at the left and right limits of the switching time $t$, denoted by $t^{-}$ and $t^{+}$,
${E(t^{+})\!-\!E(t^{-})}=\frac{k{(z(t^{+}\!)-r(t^{+}\!))}^{2}}{2}-\frac{k{(z(t^{-}\!)-r(t^{-}\!))}^{2}}{2}.$ (12)
On a period, injection of energy and dissipation must balance. An example is provided in Fig. 4, based on a switching law that captures the fundamental transitions between cutting and raising in the rocking motion:
$r$ $=r_{2}\quad\text{if}\quad\bigl(\dot{z}\geq 0\,\text{and}\,z\leq z_{2}\bigr)\;\text{or}\;\bigl(\dot{z}\leq 0\,\text{and}\,z\leq z_{1}\bigr),$
$r$ $=r_{1}\quad\text{if}\quad\bigl(\dot{z}\geq 0\,\text{and}\,z\geq z_{2}\bigr)\;\text{or}\;\bigl(\dot{z}\leq 0\,\text{and}\,z\geq z_{1}\bigr).$
The phase plane of Fig. 4 shows that solutions converge to a closed curve. Intuitively, if the mass moves too fast, the damper dissipates more energy, returning the system toward this closed curve. Conversely, if motion is too slow, less energy is dissipated, and the system gains more momentum until it approaches the same closed curve.
Figure 4: Phase plane of a mass spring damper system with switching references $r_{1}=-1$ and $r_{2}=1$ and positions $z_{1}=-0.8$ and $z_{2}=0.8$. Trajectories with different initial conditions converge to the same limit cycle.
To conclude this section, we provide a practical estimation of the robot’s work per cycle. At steady state, this metric quantifies the energetic cost of the cutting process itself (interaction with food and the environment) by isolating it from internal mechanical and virtual damping losses.
$E_{\mathcal{C}}=\int_{t}^{t+\bar{T}}\boldsymbol{F}(\tau)^{T}\dot{\boldsymbol{x}}(\tau)d\tau,$ (13)
where $\bar{T}$ is the cycle period, $\boldsymbol{F}$ is the force at the robot end-effector, and $\boldsymbol{x}$ its position. Practically this force is sampled by a force sensor, leading to the approximated formula
$E_{\mathcal{C}_{j}}\ =\!\!\!\!\sum_{t_{i}\in\mathcal{C}_{j}\setminus\{t_{N}\}}\boldsymbol{F}^{T}\bigl(t_{i}\bigr)\bigl(\boldsymbol{x}(t_{i+1})-\boldsymbol{x}(t_{i})\bigr),$ (14)
where $\mathcal{C}_{j}=\{t_{1},\dots,t_{N}\}$ contains the sampling time instants of cycle $j$ at which sample measures are taken.
Fig. 5 illustrates the work done during a carrot-cutting experiment. Initially, the knife performs a pure rocking motion without food. Between the 7th and 16th cycles, a carrot is introduced and processed, after which it is removed. Fig. 5 provide indirect evidence of the stability of the achieved periodic motion. The work done increases due to the additional dissipation of cutting through the carrot but the system returns to its baseline energy level once the carrot is removed. This behavior demonstrates the virtual model controller’s ability to compensate for external perturbations and consistently restore the original energy balance.
5 Experimental platform and cutting motion
5.1 Robot platform and control architecture
@@
The virtual model controller outlined in the previous section is implemented through the VMRobotControl.jl package Larby (2024), a platform-independent framework designed for virtual model control in robotic systems. This package enables the construction and simulation of both the robot and virtual components. Given the robot’s URDF representation, the package offers several primitives to define the virtual model controller, to compute its dynamics, and to generate the torque commands for each joint.
@@
Our main experimental platform consists of a 7-DoF Franka Research 3 arm, with a maximum payload of 3 kg. To collect force data, a ROBOTIQ FT 300-S force-torque sensor is mounted at the robot wrist, providing force and torque measurements at 100 Hz with a noise level of approximately 0.1 N. A knife is rigidly connected to the sensor using a 3D-printed attachment as shown in Fig. 1. The portability of the virtual model controller allowed us to implement and test it also on the humanoid robot Sciurus17, a lightweight and low-cost platform developed by RT Corp. Scirus17 features 17 degrees of freedom (DoF) distributed across its upper body (with 7 DoF per arm, 2 DoF for the head and 1 DoF for the torso). The robot utilizes Dynamixel motors with current control, and the conversion from torque to current is approximated by means of a linear relation based on the performance characteristics of the motor.
@@
The controller runs on a standard computer. The overall control architecture is depicted in Fig. 6. Virtual forces $\boldsymbol{f}$ generated by the virtual mechanisms are computed and translated into joint torques $\boldsymbol{\tau}$ using (1). The computation of the current state of the virtual mechanisms, such as the elongation of virtual springs and the velocities used by the virtual dampers, is based solely on the robot’s joint position $\boldsymbol{q}$ and velocity $\dot{\boldsymbol{q}}$. All force computations and their conversion into joint torques rely on forward kinematics and the associated Jacobians. This represents a significant departure from traditional motion planning, where execution is typically based on inverse kinematics.
@@
5.2 The cutting motion in experiments
@@
The virtual mechanisms constrain the motion of the blade by enforcing compliance at contact points and ‘soft’ constraints. Switching of the reference at specific points induces a cyclical motion of the knife, which is determined by the interaction among robot dynamics, compliant virtual mechanism constraints, and contacts with the environment. A first illustration is provided in Fig. 7, whose control parameters are specified in Table 1 (a). Using courgette as the cutting target, the snapshots in Fig. 7(a) illustrate the continuous contact between the knife and the cutting board while the knife rolls back and forth. The successful separation of the courgette is also seen in the snapshots. Since the motion is not pre-programmed, variations in the knife geometry or cutting board features will yield different motions.
Fig. 7(b) shows the position of two springs responsible for the rocking motion. The sliding of the virtual mass ensures that $\boldsymbol{p}_{a}$ stays close to $\boldsymbol{p}_{1}$ and no additional force is applied in the rocking direction. The displacement between $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{a}$ along the $z$ axis pulls the knife towards the cutting board and ensures contact. The knife moves to the next slice position when raised. Fig. 7(c) and 7(d) show the end-effector’s position and force, orientation and torque. The resulting end-effector’s motion is a fairly complicated trajectory, difficult to design manually. One would need to know the geometry of the knife to decide where the end-effector should be while rolling the knife back and forth.
@@
6 Experimental result and discussion
6.1 Forces and velocities